Using the financial statements and additional information, compute the following ratios for Hampton Company for 2019. Show all computations.
The financial statements of Hampton Company appear below Hampton Company Comparative Balance Sheet December 31, 2018 40,000 60,000 30,000 70,000 00,000 ASSETS 2019 ash $35,000 15,000 50,000 50,000 250,000 400,000 Short-term investments ccounts receivables nventory Property, plant & equip (net) Total assets 500,000 Time left 2:50:24

LIAB & Stockholders Equity Accounts payable Short-term notes payable Bonds payable Common stock Retained earnings 10,000 40,000 88,000 30,000 90,000 160,000 145,000 75,000 500,000 102,000 Total Liab & SE 400,000

Hampton Company Income Statement For the year ended December 31, 2019 et Sales ost of Goods Sold ross Profit xpenses Interest Expense Selling Expense Admin. Expense 360,000 162,000 12,000 40,000 59,000 Total Expense ncome before income taxes 51,000 ncome tax expense Net Income 36

Additional Information: a. Cash dividends of $9,000 were declared and paid in 2019 b. Weighted Average number of shares of common stock outstanding during 2019 was 30,000 shares. c. Market value of common stock on December 31,2019 was $21 per share. Instructions Using the financial statements and additional information, compute the following ratios for Hampton Company for 2019. Show all computations. . Current ratio: 2. Return on Common stockholders equity 3. Price-earnings ratio . Acid-test ratio 5. Receivables turnover_ 6. Tlmes interest earned 7. Profit margin

8. Days in inventory 9. Payout ratio 10. Return on assets
Answer #1
1.current ratio=current assets / current laibilties
current assets
cash 35000
short term investments 15000
accounts receivable 50000
inventory 50000
(A) 150000
Current liabilities
Accounts payable 10000
short term notes payable 40000
(B) 50000
Current ratio =A/B=           3.00
2.Return on stockholders equity
net income(A) 36000
Avg stockholders equity(B)= 241000
(160000+102000+145000+75000)/2
A/B= 14.9%
3.P/E ratio =MPS/EPS
EPS=NET INCOME / NO. OF SHARES 1.2
(36000/30000)
MPS(given) 21
P/E=21/1.2= 17.5
4.Acid test ratio
Quick assets
cash 35000
short term investments 15000
accounts receivable 50000
(A) 100000
Current liabilities
Accounts payable 10000
short term notes payable 40000
(B) 50000
Quickratio =A/B=           2.00
5. Accounts receivable turnover
Sales(A) 360000
avg. accounts receivable(B) 40000
(50000+30000)/2
AR ratio (A/B)           9.00
6.Times interest earned=EBIT/ interest expense
EBIT
net income 36000
Add:tax 15000
add:interest expense 12000
EBIT(A) 63000
interest expense(B) 12000
(A/B)           5.25
7. Profit margin
Net income(A) 36000
Sales(b) 360000
Profit margin(A/B) 10.0%
8.Days in inventory =365/ inventory turnover
COGS(a) 198000
Avg inventory(b) 60000
(50000+70000)/2
Inventory turnover (a/b)           3.30
365/inventory turnover      110.61
9.Payout ratio
Dividends(a) 9000
net income(b) 36000
Payout ratio 25.0%
10.Return on assets = net income / avg. total assets
net income(a) 36000
avg. total assets(b) 450000
(400000+50000)/2
ratio(a/b) 8.0%